# Homological stability for automorphisms of symmetric bilinear forms

### v1

@@ 4 Connectivity of the space of destabilisations
Let $x=(x_{1},...,x_{z(M)})\in\mathcal{IU}(M)$ be of maximum length, and let $(y_{1},...,y_{z(M)})$ be a witnessing sequence to it. The proposition can be proved identically to [unigrp, Lem. 6.8], by writing $M$ as $M^{\prime}\oplus e_{x,y}$ and taking $W\coloneqq\langle x_{1},...,x_{z(M)}\rangle$ instead of what they consider. To illustrate the method, let $\mathcal{P}$ be the poset in item 3. Then $\mathcal{O}(W)\cap\mathcal{P}=\mathcal{O}(W)\cap\mathcal{U}(N)_{v}$ and for any $w\in\mathcal{P}$, we have $\mathcal{O}(W)\cap\mathcal{P}_{w}=\mathcal{O}(W)\cap\mathcal{U}(N)_{vw}$. From [unigrp, Lem. 5.1] and the fact that the stable rank of a principal ideal domain is at most $2$, it follows that $\mathcal{O}(W)\cap\mathcal{P}$ and $\mathcal{O}(W)\cap\mathcal{P}_{w}$ are (z(M)3|v|)(z(M)-3-|v|)- and (z(M)3|v||w|)(z(M)-3-|v|-|w|)-connected, respectively. The conclusion now follows from [genlin, Lem. 2.13(1)]. ∎
@@
Let $\mathcal{P}$ be the poset in question. There is nothing to do if k>z(M)c(M)2k>z(M)-c(M)-2, so assume $k\leq z(M)-c(M)-2$. Lemma 3.25 yields a decomposition $\phi:M\cong N\oplus H^{z(M)-c(M)-k}$ such that $\phi(\langle v_{1},...,v_{k}\rangle)\subseteq N$ and $z(N)=c(M)+k$. Hence, there exists $T\in\text{Met}_{2}(R)$ such that $T\subseteq N$. Let $W=\phi^{-1}(H^{z(M)-c(M)-k})$. We find that $W\oplus T\subseteq M$. Now, $\mathcal{O}(W)\cap\mathcal{P}=\mathcal{O}(W)\cap\mathcal{U}^{\prime}(M)$. Let $V\coloneq\langle v_{1},...,v_{k}\rangle$; then $N$ splits as $V\oplus P$ (as modules, not as forms!) for some $P\subset N$. Given $(w_{1},...,w_{l})\in\mathcal{P}-\mathcal{O}(W)$, because $(v_{1},...,v_{k},w_{1},...,w_{l})\in\mathcal{U}(M)$, it follows that $(w_{1}-w_{1}|_{V},...,w_{l}-w_{l}|_{V})\in\mathcal{U}(M)$. Furthermore, $\mathcal{O}(W)\cap\mathcal{P}_{(w_{1},...,w_{l})}=\mathcal{O}(W)\cap\mathcal{U}^{\prime}(M)\cap\mathcal{U}(M)_{(w_{1}-w_{1}|_{V},...,w_{l}-w_{l}|_{V})}$. Propositions 4.2 and 3.20, together with [genlin, Lem. 2.13(1)], now imply that $\mathcal{P}$ is, in fact, (z(M)c(M)3k)(z(M)-c(M)-3-k)-connected. ∎
@@ 5 Applications
The remaining assertion in the statement of the theorem follows immediately from a direct examination of the arguments above. ∎
@@
As before, let us restrict ourselves to constant coefficients in our homological stability statements for the sake of brevity.
For the remainder of this section, we will assume that $R$ satisfies the conditions of Theorem 5.1. In this case, let $c(R)$ be 0 or $\text{dim}_{U(R)}(R/(2))$ depending on whether $2$ is invertible in $R$ or not. Note that $c(R)=c(M)$ for any cofinal form $M$.

### v2

@@ 4 Connectivity of the space of destabilisations
Let $x=(x_{1},...,x_{z(M)})\in\mathcal{IU}(M)$ be of maximum length, and let $(y_{1},...,y_{z(M)})$ be a witnessing sequence to it. The proposition can be proved identically to [unigrp, Lem. 6.8], by writing $M$ as $M^{\prime}\oplus e_{x,y}$ and taking $W\coloneqq\langle x_{1},...,x_{z(M)}\rangle$ instead of what they consider. To illustrate the method, let $\mathcal{P}$ be the poset in item 3. Then $\mathcal{O}(W)\cap\mathcal{P}=\mathcal{O}(W)\cap\mathcal{U}(N)_{v}$ and for any $w\in\mathcal{P}$, we have $\mathcal{O}(W)\cap\mathcal{P}_{w}=\mathcal{O}(W)\cap\mathcal{U}(N)_{vw}$. From [unigrp, Lem. 5.1] and the fact that the stable rank of a principal ideal domain is at most $2$, it follows that $\mathcal{O}(W)\cap\mathcal{P}$ and $\mathcal{O}(W)\cap\mathcal{P}_{w}$ are (z(M)3|v|)(z(M)-3-|v|)-connected and (z(M)3|v||w|)(z(M)-3-|v|-|w|)-connected, respectively. The conclusion now follows from [genlin, Lem. 2.13(1)]. ∎
@@
Let $\mathcal{P}$ be the poset in question. There is nothing to do if k>z(M)c(M)2k>z(M)-c(M)-2, so assume $k\leq z(M)-c(M)-2$. Lemma 3.27 yields a decomposition $\phi:M\cong N\oplus H^{z(M)-c(M)-k}$ such that $\phi(\langle v_{1},...,v_{k}\rangle)\subseteq N$ and $z(N)=c(M)+k$. Hence, there exists $T\in\text{Met}_{2}(R)$ such that $T\subseteq N$. Let $W=\phi^{-1}(H^{z(M)-c(M)-k})$. We find that $W\oplus T\subseteq M$. Now, $\mathcal{O}(W)\cap\mathcal{P}=\mathcal{O}(W)\cap\mathcal{U}^{\prime}(M)$. Let $V\coloneq\langle v_{1},...,v_{k}\rangle$; then $N$ splits as $V\oplus P$ (as modules, not as forms!) for some $P\subset N$. Given $(w_{1},...,w_{l})\in\mathcal{P}-\mathcal{O}(W)$, because $(v_{1},...,v_{k},w_{1},...,w_{l})\in\mathcal{U}(M)$, it follows that $(w_{1}-w_{1}|_{V},...,w_{l}-w_{l}|_{V})\in\mathcal{U}(M)$. Furthermore, $\mathcal{O}(W)\cap\mathcal{P}_{(w_{1},...,w_{l})}=\mathcal{O}(W)\cap\mathcal{U}^{\prime}(M)\cap\mathcal{U}(M)_{(w_{1}-w_{1}|_{V},...,w_{l}-w_{l}|_{V})}$. Propositions 4.2 and 3.22, together with [genlin, Lem. 2.13(1)], now imply that $\mathcal{P}$ is, in fact, (z(M)c(M)3k)(z(M)-c(M)-3-k)-connected. ∎
@@ 5 Applications
The remaining assertion in the statement of the theorem follows immediately from an examination of the arguments above. ∎
@@
Of independent interest is that all number rings, not just the ones satisfying Assumption 3.4, admit cofinal forms:
Proposition 5.3.
For any number ring $S$, there is a cofinal form in $\textup{Unimod}^{s}(S)^{\simeq}$.
Proof.
It is enough to find a metabolic form that is cofinal among all metabolic forms. The ring $S/(2)$ is finite because $S$ is a lattice over $\mathbb{Z}$; therefore the form $\bigoplus_{s\in S/(2)}\theta(s)$ is cofinal among all metabolic forms with free Lagrangian, as follows from Propositions 3.8 and 3.9. It now suffices to show that such metabolic forms are cofinal in all metabolic forms. Let $M$ be a metabolic form with Lagrangian $L$; because $S$ is a Dedekind domain, we find that $L$ is isomorphic to $S^{n}\oplus\mathcal{L}$ for some line bundle $\mathcal{L}$. As the Picard group $\text{Pic}(S)$ is finite, we can set $t$ to be the order of $\mathcal{L}\in\text{Pic}(S)$. The form $M^{t}$ then admits as Lagrangian the module $S^{nt}\oplus\mathcal{L}^{t}\cong S^{nt+t-1}\oplus\mathcal{L}^{\otimes t}\cong S^{(n+1)t}$, which is free. This concludes the proof. ∎
For the remainder of this section, we will assume that $R$ satisfies the conditions of Theorem 5.1. In this case, let $c(R)$ be 0 or $\text{dim}_{U(R)}(R/(2))$ depending on whether $2$ is invertible in $R$ or not. Note that $c(R)=c(M)$ for any cofinal form $M$. We restrict ourselves to constant coefficients in our homological stability statements for the sake of brevity.